A grocery store has 12 lbs of a mixture of nuts and dried fruit that is 10% banana chips. If they mix additional banana chips to bring the new mixture up to 25%, how many pounds of banana chips should they add?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

percent ---------------- Amount
Mixture 10 ---------------- 12
Banana chips 100 ---------------- x
Total 25 ---------------- 12 + x

10*12+100x=25(12+x)
120+100 x=300+25x
100x-25x=-120+300
75x=180
/75
x=2.4
2.4lb of Banana chips

Answer by stanbon(75887)

A grocery store has 12 lbs of a mixture of nuts and dried fruit that is 10% banana chips. If they mix additional banana chips to bring the new mixture up to 25%, how many pounds of banana chips should they add?
--------------------------
Equation:
chips + chips = chips
0.10*12 + x = 0.25(12+x)
Multiply thru by 100 to get:
10*12 + 100x = 25*12 + 25x
-----
75x = 15*12
x = (1/5)12
x = 2.4 lbs (amt. of banana chips to be added)
===========================================
Cheers,
Stan H.